1. An RL circuit has emf of 24t volts, a resistance of 10 k-ohms, an inductance of 5H with an initial current of 1 mA. What is the current after 0.05 seconds? Ans. 118.8µA
  2. Initially a tank holds 75 gal of a brine containing 12 lb of salt. At t=0, fresh water is poured into the tank at the rate of 3 gal/min, while the well stirred mixture leaves the tank at the rate of 2 gal/min. what is the amount of salt in the tank after 1 hour? Ans. 3.70 lb
1
Expert's answer
2021-11-19T06:46:50-0500

1.


"Li'+Ri=V""5i'+10\cdot10^{3}i=24t"


"5i'=24t-10000i"

Substitute


"u=24t-10000i"


"u'=24-10000i'"

"i'=0.0024-0.0001u'"

"0.0005(24-u')=u"

"u'=-2000(u-0.012)"


"\dfrac{du}{u-0.012}=-2000dt"

Integrate


"u=0.012+c_1e^{-2000t}"


"24t-10000i=0.012+c_1e^{-2000t}"

"i=c_2e^{-2000t}+0.0024t-0.0000012"

Given "i(0)=10^{-3}\ A"

"0.001=c_2-0.0000012"


"c_2=-0.0009988e^{-2000t}+0.0024t-0.0000012"


"i(t)=-0.0009988e^{-2000t}+0.0024t-0.0000012"




"i(0.05)=-0.0009988e^{-2000(0.05)}+0.0024(0.05)"


"-0.0000012=0.0001188(A)"

"118.8\ \mu A"


2.

Let "s(t) =" amount, in lb of salt at time "t." Then we have

"\dfrac{ds}{dt}="(rate of salt into tank) − (rate of salt out of tank)


"\dfrac{ds}{dt}=0-\dfrac{2s}{75+(3-2)t}"

So we get the differential equation


"\dfrac{ds}{dt}=-\dfrac{2s}{75+t}"


"\int \dfrac{ds}{s}=-\int\dfrac{2dt}{75+t}"

"s(t)=c_1(75+t)^{-2}"

"s(0)=c_1(75)^{-2}=12"

"c_1=67500"

"s(t)=67500(75+t)^{-2}"

"s(60)=67500(75+60)^{-2}"

"s(60)=3.7037\ lb"


"3.70\ lb"